OpenStudy (anonymous):

prove the identity. show all work. tan(x-pi/4)=tanx-1/tanx+1

hartnn (hartnn):

know tan (A-B) formula ?

OpenStudy (anonymous):

tanx-tany/1+tanxtany

hartnn (hartnn):

\(\large tan(A-B)= \frac{tan A-tan B}{1+tan Atan B}\) what would u put A and B as ?

hartnn (hartnn):

thats correct, put x=? , y= ?

OpenStudy (anonymous):

a=x b=pi/4

hartnn (hartnn):

correct, what is tan pi/4= ?

OpenStudy (anonymous):

1

hartnn (hartnn):

now just plug in values...

OpenStudy (anonymous):

got it !(:

hartnn (hartnn):

good :)